OpenStudy (anonymous):

Two positive numbers have a difference of 4 and a product of 96. What are the numbers? a. 12 and 8 b. 16 and 6 c. 16 and 12 d. 18 and 14 Help me please.

OpenStudy (dan815):

y-x=4 y*x=96 solve for x and y

OpenStudy (dan815):

lets rearrange the top equation y-x=4 y=x+4 now we can sub this into the 2nd equation y*x=96 (x+4)*x=96 x^2+4x=96 solve for x

OpenStudy (anonymous):

o.o

OpenStudy (dan815):

x^2+4x-96=0 (x+12)(x-8)=0 x=-12 or 8 but we only want position so x=8 so y = 12 since y=x+4 =8+4

OpenStudy (dan815):

you could have just looking at the answers too.. and done it without the algebra like, look at each choice, see which one has a difference of 4 and multiplies to 96

OpenStudy (dan815):

could have done*

OpenStudy (anonymous):

okay A is correct.